OpenStudy (anonymous):

Trigonometry cot squared x+cscx-4 is equivalent to a.)csc squared x+csc x-5 b)csc squared 2x+cscx-3 c)cos^2x/sin^2 +1/cosx -4 d)-4 How is this question done?I need help

OpenStudy (anonymous):

i believe it is C, because \[\cot x = \frac{ \cos x }{ \sin x }\] and \[\csc x = \frac{ 1 }{ \cos x }\]

OpenStudy (anonymous):

Okk thanks but u i still dnt understand

OpenStudy (luigi0210):

actually that's wrong..

OpenStudy (anonymous):

how is it suppose to be done?

OpenStudy (luigi0210):

substitute cot^2x with csc^2x-1

OpenStudy (anonymous):

ok

OpenStudy (luigi0210):

ditto

OpenStudy (luigi0210):

ditto=same here;agreed? ditto as in i agree with your answer choice as well

OpenStudy (loser66):

so, you do it

OpenStudy (luigi0210):